# Ultrastable nanophotonic microcavities via integrated thermometry

### v1

@@ III Fully-Integrated Photonic Frequency Reference
Figure 2(c) shows the numerically-computed Allan deviation (ADEV). As expected, the stabilized cavity offers a significantly lower ADEV compared to the free-running and open-loop cavities. The shape of the ADEV of the stabilized cavity closely matches that of pure $1/f$ noise, indicating that unbounded random walk resonance frequency drifts arising from environmental and crosstalk perturbations are effectively canceled out by the stabilization method employed and that the residual fluctuations are bounded, limited only by noise in the control electronics used in this demonstration. This is evidenced by the flat shape of the ADEV versus averaging time, compared to the free-running and open-loop cavities, which exhibit non-stationary unbounded random walks, shown by the increasing ADEV at large averaging times. The stabilized microresonator also offers significantly higher long-term stability compared to temperature-stabilized DFB lasers, which are commonly used in chip-based telecommunications applications [49, 30], making it a highly practical fully-integrated frequency reference cavity.
Finally, long-term repeatability of the microcavity’s tunable absolute resonance wavelength is assessed as follows. To assess the long-term repeatability of a thermally-tuned cavity only subject to ambient thermal noise (i.e., $\xi=0$), the voltage applied to the outer heater is swept in discrete steps so as to trace the logo of Columbia University, and the microcavity’s absolute resonance wavelength is synchronously measured using the scanning probe laser setup described in Fig. 2(a). This measurement is repeated in the presence of additional perturbations to mimic external thermal fluctuations in practical photonic devices ($\xi=3.77$). No feedback is applied in the first two configurations. Finally, the feedback loop is closed, and thermal tuning is performed by simply changing the set-point of the PID loop. The results are shown in Fig. 3.
Figure 3: Demonstration of tunability and long-term repeatability of the microresonator’s absolute resonance wavelength. The resonance frequency is programmed to trace a pre-defined shape by blind tuning of the power supplied to the outer heater in the presence of ambient noise alone ($\xi$=0; shown in a), in the presence of additional perturbation ($\xi$=3.77; shown in b), and by tuning of the set-point for temperature stabilization in the presence of ambient noise as well as additional perturbation ($\xi$=3.77; shown in c). The measurements for heater-only tuning show that the microresonator’s resonance wavelength exhibits a large drift over the timescale of only a few hours, even for the same pre-programmed heater voltage. The stabilized tuning, on the other hand, yields a highly repeatable performance even in the presence of strong ambient and external noise.
When tuned with the heater alone, i.e., the most common way of tuning microresonators, applying the same voltage does not lead to the same resonance wavelength over the measurement duration of 10 hours, as evidenced by the drift in the patterns in Figs. 3(a) and (b). Similar to the results presented in Fig. 2, additional perturbations lead to “short-term” fluctuations superimposed on top of the long-term drift caused by environmental drift. However, in the stabilized case in which the set-point of the feedback loop is tuned, active feedback based on the resistance thermometer’s “reading” cancels out ambient drift as well as the perturbation-induced fluctuations, and the microcavity resonance wavelength exhibits significantly better repeatability.
The long-term frequency stability of the thermometrically stabilized microresonator is demonstrated further by stabilizing the frequency of a laser. A DFB laser is locked to a microcavity resonance using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is recorded every 10 seconds using a bench-top wavelength meter. We employ the PDH method in this experiment due to its insensitivity to drift in the input fiber-to-chip optical coupling. In a fully-integrated system, however, the frequency stability of the microresonator is easily transferred to an integrated laser via mechanisms such as self-injection locking [50]. The results of the frequency stabilization demonstration are shown in Fig. 4.
Figure 4(a) shows the measured drift in laser center wavelength for the free-running and PDH-locked DFB lasers, in blue and maroon respectively. The mean emission wavelength for the two cases are approximately 1551.2 nm. The free-running DFB experiences significant wavelength drift arising from ambient temperature fluctuations alone, amounting to a standard deviation of $11.5$ pm (i.e., $1.43$ GHz) over the span of $42$ hours of measurement time. In comparison, the locked DFB emits at a highly stable center wavelength, with a peak-to-peak variation that lies within ±$\pm 0.5$ pm of the mean over $50$ hours of measurement time, and exhibits a standard deviation of $0.24$ pm. The stability enhancement factor is 48$\approx$ 48$\times$ 48 ×. The long-term stability measured here is significantly superior to many commercially-available DFB laser system, and makes the scheme presented here ideal for the realization of practical photonic transceivers for dense wavelength division multiplexing (DWDM) applications.
Figure 4(b) shows that the computed ADEV of the locked DFB is significantly lower ADEV than the free-running DFB, and is comparable to the ADEV of the stabilized microresonator displayed in Fig. 2. By using a large integral gain in the feedback servo of the PDH, a slightly lower ADEV is observed at smaller averaging times. The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
Figure 4: Demonstration of long-term laser frequency stabilization to the microresonator. A DFB laser is locked to the resonance of the stabilized microresonator using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is continually measured once every 10 seconds using a bench-top wavelength meter. a, Time-domain comparison of center wavelength drift of free-running and PDH-locked configurations of a DFB laser. The locked laser’s center wavelength remains within a ±$\pm 0.5$ pm window of the mean, indicated by the dashed lines, for 50 hours. The free-running laser is only subject to ambient temperature drift in a temperature-stabilized laboratory environment. b, Comparison of measured Allan Deviation (ADEV) of the free-running and locked configurations of the DFB laser. The locked laser exhibits slightly lower ADEV than the stabilized microresonator at lower averaging times due to the strong integral gain of the PDH feedback servo. The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
IV Fully-thermally-stabilized Kerr Comb
@@
Finally, we show that the fully-integrated thermometry approach described in this work can stabilize a modelocked Kerr comb, which is typically highly sensitive to perturbations, for long periods. The experimental schematic is shown in Fig. 5(a). A pump laser at approximately 1550 nm is amplified before coupling into the microcavity device. The off-resonance optical power output is 150-mW. The low-noise current source connected to the resistance thermometer is modulated with a burst waveform to generate a single-soliton Kerr comb, shown in blue in the optical spectrum analyzer (OSA) measurements in Fig. 5(b). A portion of the out-coupled Kerr comb is sent through a fiber-optic notch filter centered at the pump wavelength, and the transmitted output, shown in red in Figure 5(b), is used to track the comb power using a photodiode. Phase-noise characterization is performed by generating a microwave signal at the repetition rate of the comb (76 GHz) via a high-speed photodetector (bandwidth 100 GHz), and measuring the microwave phase noise using a phase noise analyzer (PNA; Rohde & Schwarz ®  FSWP50). Since the microwave frequency (76 GHz) is beyond the specified bandwidth (50 GHz) of the PNA used in this work, an 8:1 harmonic mixer (Rohde & Schwarz ®  RPG FS-Z110) is employed to down-convert the microwave frequency to within the PNA’s bandwidth.
@@
The stability of the Kerr comb is studied by applying a perturbation (white-noise voltage, bandwidth $=20$ Hz) via the outer on-chip heater and measuring the real-time on-chip temperature and the power and phase noise of the comb. For a noise strength of $\xi=1.51$, without any feedback correction, the modelocked Kerr comb transitions to a chaotic comb state almost immediately, as shown in Fig. 5(c), where the top plot shows the applied thermal perturbation voltage signal in the time-domain. The thermometer voltage plot, for a constant low-noise current sent through the resistance thermometer, shows a sudden increase in the cavity temperature, indicating a sudden jump to a state with higher intracavity power. This is explained by the bottom plot, which shows that the comb power (which is measured by filtering out the pump line) is no longer stable and exhibits random fluctuations driven by the perturbation voltage. These fluctuations are also seen in the thermometer voltage following the jump. The smaller amplitude of fluctuations observed by the thermometer is a result of an increase (decrease) in temperature at the thermometer due the perturbation and the consequent secondary effect, namely decrease (increase) in temperature at the thermometer due to decrease (increase) in intracavity optical power and self-heating. This is a result of the inverse relationship between cavity temperature and intracavity power for the effectively blue-detuned pump of the chaotic comb state, which does not apply for the modelocked comb state, since the pump is effectively red-detuned, and the resistance thermometer therefore remains sensitive to external perturbations, making it an efficient tool in stabilizing modelocked comb states.
@@
The stability of the closed-loop system is studied by repeating the same measurement sequence as above. However, in this case, the noise strength is increased five-fold to $\xi=7.55$. As seen in Fig. 5(d), the PID correction voltage cancels out the applied perturbation, leading to a stable temperature and Kerr comb power even in the presence of strong perturbations. Whereas the Kerr comb is pushed out of its existence range immediately in the open-loop ($\xi=1.51$) scenario, closed-loop operation was observed to support modelocking for over 20 minutes, limited only by the drift in the input fiber-to-chip optical coupling, paving the way for Kerr-comb-based photonic devices that can operate in the desired modelocked state indefinitely.
Lastly, Fig. 5(e) shows the comparison of the microwave single-sideband (SSB) frequency noise. For the ‘open-loop’ scenario, the noise strength is reduced to $\xi=0.75$ to keep the comb in the modelocked regime of operation. For the ‘stabilized’ scenario, the noise strength is maintained at $\xi=7.55$. As the yellow and purple solid curves show, the frequency noise of the microwave generated by the stabilized cavity is over 20 dB below that of the microwave generated by the open-loop cavity. At lower offset frequencies (i.e., over longer timescales), the stabilized cavity’s microwave exhibits a lower noise than the free-running cavity’s microwave since ambient drift is more prominent on longer timescales.
V Conclusion
In conclusion, we demonstrate a powerful new approach based on fully-integrated thermometry to stabilize the absolute resonance frequency of high-$Q$ silicon nitride microcavities without requiring more complicated monitoring systems such as photodetection. The thermometric stabilization mitigates environmental drift as well as externally-induced perturbations effectively, and the Allan deviation of the stabilized microcavity’s resonance frequency remains nearly flat versus averaging time, in contrast to the unbounded random walks exhibited by the free-running and open-loop cases, evidenced by the increasing Allan deviation at large averaging times. By locking a DFB laser to the stabilized microresonator, we demonstrate that the center-wavelength stability of the laser is <±$<\pm 0.5$ pm from the mean over the measurement duration of $50$ continuous hours, which is a 48×48$\times$48 × improvement compared to the free-running laser. This work paves the way for highly compact, fully-integrated and field-deployable frequency references for telecommunications applications.
When pumped with a high-power laser source to generate a Kerr comb, the thermometry-based stabilization technique described in this work makes the comb significantly more stable even in the presence of strong perturbations. The fully-thermally-stabilized comb demonstrated here offers a path toward comb-based high-$Q$ photonic devices that can operate indefinitely even in noisy environments.
Although the demonstration in this work was limited to the SiN platform, this technique is readily generalizable to other photonic materials. Notably, previously-demonstrated on-chip temperature-sensing techniques, such as ones leveraging the photoconductive effect, require the presence of significant optical power in the cavity, and are limited to semiconductor platforms. The resistance-thermometer-based scheme here poses no such requirement. Finally, removing the requirement of additional drop-ports and photodetection for stabilization can reduce the design complexity and fabrication costs of practical photonic devices quite significantly.
VI Acknowledgments
VII Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
VIII Funding
The authors acknowledge support from the Army Research Office under Cooperative Agreement Number W911NF-24-2-0204, the Air Force Office of Scientific Research (FA9550-20-1-0297), and the Defense Advanced Research Projects Agency (HR0011-19-2-0014).
Ultrastable nanophotonic microcavities via integrated thermometry: Methods
I Thermal Design for Integrated Thermometry
Figure M1: Thermal design for integrated thermometry. a, Numerically simulated steady-state temperature profile for the scheme described in this work. The heater power is set to $100$ mW for the purposes of illustration. The initial temperature of the chip is $293$ K. b, Heater-thermometer spacing design trade-off, showing the core-thermometer temperature differential (left y-axis, in blue), and the heater power required for core heating of $0.1$ K (right y-axis, in dark red), versus heater-thermometer spacing.

### v2

@@ I Results
Figure 2(c) shows the Allan deviation (ADEV) versus averaging time computed from the time-domain data. The stabilized cavity offers a significantly lower ADEV compared to the free-running and open-loop cases. The flat shape of the stabilized cavity’s ADEV closely matches that of pure $1/f$ noise, indicating that unbounded random-walk resonance frequency drifts arising from environmental and crosstalk perturbations are effectively canceled out by the stabilization method, and that residual fluctuations are bounded and limited only by noise in the control electronics used in this work. Supplementary Fig. S4 provides ADEV comparison measurements demonstrating that the stabilized microresonator offers superior long-term frequency stability compared to many commercially available telecommunication lasers [56, 37] available in our laboratory, demonstrating its potential utility as a highly practical integrated reference cavity.
The long-term repeatability of the microcavity’s tunable resonance wavelength is illustrated by the demonstration of a resonance-wavelength thermometric lookup table. A one-time calibration measurement of the absolute resonance wavelength versus the thermometer resistance is performed using the setup described in Fig. 3a. The current source output is monitored by measuring the voltage across a series $50.0$ $\Omega$ bulk termination resistor using a precision 6.5-digit digital multimeter (DMM; Agilent 34401A). The thermometer voltage is measured by subtracting the voltage across the termination from the total voltage across the current source, which in stabilized operation equals the PID set-point voltage. The thermometer resistance calculated as the ratio of these two quantities also includes a small contribution from contact resistance of the spring-loaded electrical contact probes (PicoProbe 40A-GS-150-P) employed in our experimental setup. The results of the one-time calibration are shown in Fig. 3b. The linear fit extracted from this calibration is used in subsequent measurements to compute the desired thermometer resistance to achieve any desired resonance wavelength within the calibration range.
A computer-generated string of $10,000$ pseudorandom “target” resonance wavelengths is fed into a thermometric control loop implemented using a field-programmable gate array (FPGA), which calculates the desired thermometer resistance corresponding to a target resonance wavelength using the calibration fit, and stabilizes the microresonator to the computed temperature. The resonance wavelength is then measured and compared with the target value. Each such “measurement” takes approximately five seconds. Two separate sets of 10,000 measurements each are performed. A zoomed-in portion of the first set of 10,000 measurements, which concluded 12 hours after the calibration and is displayed in Fig. 3c, shows close agreement between the measured and target resonance wavelengths. Figure 3(d) shows histograms of the absolute-wavelength error (bin size =0.1=0.1 pm), defined as the target minus measured wavelengths. For the first set, a root-mean squared error (RMSE) is $0.11$ pm is observed. A systematic error, resulting in a larger RMSE of $0.77$ pm, was observed in the second set carried out 3 days after the calibration. It is hypothesized to be stemming from drift in the contact resistance of the spring-loaded probes.
The repeatability of the thermometrically controlled microcavity’s resonance wavelength is further demonstrated by programming the resonance wavelength to trace a predetermined pattern. When the heater is actuated without feedback control, in the presence ($\xi=3.77$) as well as absence ($\xi=0$) of additional perturbation, significant long-term drift is observed over the $10$-hour measurement duration (Figs. 4(a) and (b)). On the other hand, when the microresonator is stabilized and the set-point voltage is tuned instead, long-term drift is canceled out, and the resonance wavelength exhibits a highly repeatable relationship with the set-point voltage, as shown in Fig. 4(c).
The frequency stability of the microresonator is transferred to a fiber-coupled DFB laser using the Pound-Drever-Hall (PDH) method. For an integrated laser scheme, self-injection locking may be more suitable [22]. The DFB emission wavelength is recorded once every 10 seconds using a wavelength meter. As shown in Fig. 5(a), the free-running DFB experiences significant wavelength drift due to ambient temperature fluctuations, with a standard deviation of $11.5$ pm (i.e., $1.43$ GHz) over $42$ hours of measurement time. In comparison, the locked DFB is highly stable, with a peak-to-peak variation in emission wavelength that $\pm 0.5$ pm of the mean for over $50$ hours, and a standard deviation of $0.24$ pm (i.e., $30$ MHz). The stability enhancement factor is $\approx 48\times$. The long-term stability reported here is superior to many commercially available DFB and wavelength-locker-based laser systems. Figure 5(b) shows that the ADEV of the locked DFB is significantly lower and is comparable to the ADEV of the stabilized microresonator displayed in Fig. 2. Due to the large integral gain of the PDH servo, a lower ADEV is observed at smaller averaging times. The higher ADEV at larger averaging times is attributed to insufficient gain of the PDH servo at very low frequencies.
@@
Finally, we show in Fig. 6(a) an approach to stabilize a modelocked Kerr comb, which is typically highly sensitive to thermal perturbations. A $1550$-nm pump laser is amplified before coupling into the photonic chip (pump power =150=150 mW). The thermometer current source is modulated to generate a single-soliton Kerr comb [49]. The comb power is measured using a photodiode after filtering the pump line using a fiber-optic notch filter. The stability of the Kerr comb is studied by applying a white-noise perturbation (bandwidth =20=20 Hz) via the outer on-chip heater and measuring the on-chip temperature and the output comb power. The repetition-rate noise is measured by generating a microwave using a $100$-GHz bandwidth photodetector and characterizing it using a phase noise analyzer (PNA; Rohde & Schwarz ®  FSWP50) via an 8:1 harmonic mixer.
@@
For $\xi=1.51$ and without feedback correction, the modelocked Kerr comb transitions to a chaotic comb state almost immediately, as seen in Fig. 6(c). The thermometer voltage plot shows a sudden increase, indicating an abrupt jump to a state with higher intracavity power, such as a chaotic comb. This is corroborated by the random fluctuations exhibited by the comb power and the thermometer voltage following the jump. The smaller amplitude of fluctuations observed by the thermometer following the jump result from an increase (decrease) in temperature at the thermometer due the perturbation and its consequent secondary effect, namely decrease (increase) in thermometer temperature due to decrease (increase) in intracavity optical power and self-heating. This is caused by the effectively blue-detuned pump of the chaotic comb and is not applicable to the modelocked comb due to its effectively red-detuned pump.
In the stabilized case ($\xi=7.55$; Fig. 6(d)), the PID correction voltage cancels the applied perturbation, leading to a stable temperature and comb power. Whereas the Kerr comb is pushed out of its existence range immediately in the open-loop ($\xi=1.51$) scenario, closed-loop operation was observed to support modelocking for over 20 minutes, limited only by the drift in the input fiber-to-chip optical coupling. In the absence of coupling drift, such as in packaged photonic devices, cold-cavity detuning can be made impervious to external thermal fluctuations. Thermometric control can then enable Kerr-comb-based photonic devices that potentially remain modelocked indefinitely.
Lastly, Fig. 6(e) shows the measured microwave single-sideband (SSB) frequency noise spectra. For the ‘open-loop’ scenario, the noise strength is reduced ($\xi=0.75$) to maintain modelocking. The microwave noise for the stabilized case is >20>20 dB below the open-loop case despite the stronger added-noise strength ($\xi=7.55$). At offset frequencies below $\sim 1$ Hz, the stabilized cavity’s microwave exhibits lower noise than the free-running microwave since ambient drift is more prominent on longer timescales. A comparison of the measured soliton repetition rates is presented in Supplementary Fig. S5.
II Discussion
In summary, we have demonstrated a powerful new approach based on fully integrated thermometry to stabilize the resonance frequency of high-$Q$ microresonators without requiring any photodetection. Thermometric stabilization mitigates environmental as well as crosstalk-induced thermal perturbations, delivering a highly stable microresonator whose residual frequency fluctuations are only limited by $1/f$ noise in the control electronics. Realizing temperature-stabilized microresonators that are TRN-limited on long timescales is an exciting topic for future research, as it offers the prospect of precision metrology on a chip. We demonstrate that after a one-time calibration, the absolute resonance wavelength can be tuned to any desired value within the calibration range via thermometric control, with measured absolute-wavelength RMSEs $0.11$ pm and <0.77<0.77 pm after $12$ hours and $3$ days following the calibration, respectively. The emission wavelength of a DFB laser frequency-locked to the microresonator remains within $\pm 0.5$ pm of the mean over $50$ continuous hours, which is a $48\times$ improvement compared to free-running operation. This performance is superior to many commercial DFB and wavelength-locker-based laser systems. Finally, we stabilize a Kerr comb against strong thermal perturbations, offering a path toward chip-based comb devices that can potentially remain modelocked indefinitely in noisy environments.
Although our demonstration was focused on SiN microresonators, our technique is readily applicable to other dielectric and semiconductor photonic materials. Removing the requirement of drop-ports and photodetection for stabilization can significantly reduce design complexity and fabrication costs of practical photonic devices. Finally, the concepts described here can be applied to several other photonic integrated systems, including microresonator arrays and Mach-Zehnder interferometer (MZI) meshes for topological photonics [57] and photonic computing [58, 59].
III Acknowledgments
The authors acknowledge support from the Army Research Office under Cooperative Agreement Number W911NF-24-2-0204 (M.L. and A.L.G.), the Air Force Office of Scientific Research under FA9550-20-1-0297 (M.L. and A.L.G.), and the Defense Advanced Research Projects Agency under HR0011-19-2-0014 (M.L. and A.L.G.).
IV Author Contributions
S.K.D. and A.L.G. conceived the project. S.K.D. and Y.Z. performed the experiments. S.K.D. and G.R.B. performed the design and characterization of the resistance thermometer. S.K.D., Y.Z., and A.L.G. performed the data analysis with input from all authors. K.J.M. fabricated the devices under the supervision of M.L. S.K.D. and A.L.G. wrote the manuscript with feedback from all authors. M.L. and A.L.G. supervised the project.
V Competing Interests
S.K.D., Y.Z., K.J.M., G.R.B., M.L., and A.L.G. are inventors on a U.S. provisional patent application filed by The Trustees of Columbia University in the City of New York (application number 63/830,566) on 26 June 2025 that covers ultrastable microresonators using integrated thin-film resistance thermometers and thermistors.
Figure 1: Integrated thermometry. a, Conceptual illustration of integrated thermometry applied to stabilize a high-Q monolithic microresonator against thermal fluctuations arising from ambient heat sources and crosstalk from other thermally-tuned devices on the same chip. A thin-film metallic (Platinum) resistor, typically used as a microheater, exhibits a temperature-dependent resistance due to the metal’s natural temperature-dependent resistivity. As a result of the low heat capacity of the thin-film resistor, small fluctuations in the heat flow in its vicinity lead to large changes in its temperature, enabling the real-time monitoring of on-chip temperature by simply measuring its electrical resistance. A second identical resistor is used as a heater to perform active stabilization purely using the resistance thermometer, thereby eliminating the need for optical probing for stabilization. Image for data center sourced from Adobe Stock, Ref. [61]. Image courtesy for picture of co-packaged device: Kaylx Jang. b, Current-voltage-resistance (I-V-R) characteristics of the resistance thermometer fabricated in this work. The I-V measurements (pale blue dots) deviate from the linear trend (dark blue line), and the resistance (yellow dots) exhibits a quadratic dependence (purple line) on the thermometer voltage. c, The measured thermometer resistance (lavender dots) also exhibits a quadratic dependence (green line) on the voltage applied across the (second) heater, indicating a linear dependence on power dissipated across the heater. d, Resonance frequency shift (green circles) of a high-Q microcavity of free-spectral range (FSR) 76 GHz, measured using a calibrated piezo-tuned probe laser, and measured change in thermometer resistance (gold line), for a sinusoidal perturbation applied via the heater. The cavity temperature and resonance frequency exhibit a strong anti-correlation. e, Measured resonance frequency shift plotted versus temperature measured by the thermometer, same data as in d. The linear fit provides an in-situ measurement of the thermo-optic coefficient of the fundamental spatial mode of the cavity waveguide.
Figure 2: Long-term stabilization of the absolute resonance frequency of a high-Q SiN microresonator using integrated thermometry. a, Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance. Absolute frequency calibration of the scan is performed by producing a heterodyne note between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength drift using a benchtop wavelength meter. The heterodyne beat note as well as the cavity resonance line-shape are measured synchronously with the piezo sweep using a real-time oscilloscope connected to a computer interface for periodic data acquisition every 10 seconds. For $\xi$=1, the pseudorandom perturbation applied via the heater yields a root mean-squared (RMS) of the induced cavity resonance frequency fluctuation that is equal to the cavity resonance full-width-at-half-maximum (FWHM) linewidth, which is 75 MHz for the microcavity fabricated in this work. b, Drift in the microcavity’s calibrated resonance frequency over the duration of 24 hours, for free-running ($\xi$=0), open-loop ($\xi$=7.55), and stabilized ($\xi$=7.55) operation. The free-running cavity exhibits a slow but strong drift in its resonance frequency due to ambient temperature drift in the lab, whereas the open-loop cavity additionally exhibits strong “fast” fluctuations due to the perturbation introduced. The stabilized cavity’s resonance frequency remains highly stable in the presence of very strong ambient and crosstalk-induced perturbations. c, Allan Deviation (ADEV) versus averaging time, computed using the data displayed in b. The stabilized cavity exhibits an ADEV that is well below the free-running and open-loop modes of operation. The shape of the ADEV curve of the stabilized cavity matches that of $1/f$ noise, indicating that performance is only limited by control electronics. d, Histogram plots depicting the fraction of time over 24 hours (y-axis) that the microcavity’s absolute resonance frequency drift lies within a given frequency bin with a width of 5 MHz.
Figure 3: Demonstration of thermometric absolute-resonance-wavelength lookup table. a., Simplified experimental schematic, depicting a tunable probe laser that is piezo-scanned across a microcavity resonance. The absolute resonance wavelength is measured by producing a heterodyne beat between the tunable laser and a stable reference laser, and simultaneously tracking the reference laser’s wavelength using a highly-stable benchtop wavelength meter. The current produced by the low-noise current source is monitored by measuring the voltage across a series $50.0$ $\Omega$ bulk resistor using a precision 6.5-digit digital multimeter (DMM). The PID loop is implemented using a commercial FPGA, and the set-point voltage is input via a computer interface. b. Calibration measurement of the absolute resonance wavelength vs the thermometer resistance. The linear fit extracted from this one-time calibration is utilized in subsequent measurements to compute the desired thermometer resistance to achieve any desired resonance wavelength within the calibration range. c. Demonstration of arbitrary absolute-wavelength tuning of the microresonator via thermometry. A string of 10,000 pseudorandom “target” wavelengths (red hollow circles) are provided as inputs to the thermometric control system illustrated in a., which computes the desired temperature for each target wavelength and stabilizes the microresonator to the computed temperature. The resonance wavelength is measured (blue-magenta filled circles) and compared with the target wavelengths. A zoomed-in portion of 100 consecutive measurements out of 10,000 is shown. d. Histograms showing the deviation between the target and measured absolute resonance wavelengths following a one-time calibration. The histogram bin size is 0.1 pm. The first set of 10,000 measurements, which concluded 12 hours after the calibration run, exhibits a root-mean squared absolute-wavelength error of 0.11 pm. The second set of measurements three days later shows a similar narrow distribution of absolute-wavelength error, but exhibits a systematic error, resulting in a 0.77 pm.
Figure 4: Tunability and long-term repeatability of the microresonator’s absolute resonance wavelength. The resonance frequency is programmed to trace a pre-defined shape by blind tuning of the power supplied to the outer heater in the presence of ambient noise alone ($\xi$=0; shown in a), in the presence of additional perturbation ($\xi$=3.77; shown in b), and by tuning of the set-point for temperature stabilization in the presence of ambient noise as well as additional perturbation ($\xi$=3.77; shown in c). The measurements for heater-only tuning show that the microresonator’s resonance wavelength exhibits a large drift over the timescale of only a few hours, even for the same pre-programmed heater voltage. The stabilized tuning, on the other hand, yields a highly repeatable performance even in the presence of strong ambient and external noise.
Figure 5: Demonstration of long-term laser frequency stabilization to the microresonator. A DFB laser is locked to the resonance of the stabilized microresonator using the Pound-Drever-Hall (PDH) method, and its emission center wavelength is continually measured once every 10 seconds using a bench-top wavelength meter. a, Time-domain comparison of center wavelength drift of free-running and PDH-locked configurations of a DFB laser. The locked laser’s center wavelength remains within a $\pm 0.5$ pm window of the mean, indicated by the dashed lines, for 50 hours. The free-running laser is only subject to ambient temperature drift in a temperature-stabilized laboratory environment. b, Comparison of measured Allan Deviation (ADEV) of the free-running and locked configurations of the DFB laser. The locked laser exhibits slightly lower ADEV than the stabilized microresonator at lower averaging times due to the strong integral gain of the PDH feedback servo. The slightly higher ADEV at larger averaging times is attributed to insufficient gain of the PDH feedback servo at very low frequencies.
@@ V Competing Interests
Frequency-stable nanophotonic microcavities via integrated thermometry: METHODS
I Device Fabrication
II Calibrated Measurement of Absolute Resonance Frequency
[…]
III Thermal Design for Integrated Thermometry